Hello, I am in Algebra II.
There are two positive numbers with the following properties.
1. The first number is 4 more than the second number.
2. The product of the two numbers is 4 more than their sum.
What could these two numbers be?

Answer by mananth(16946)

one number be x
Other number will be x+4
The product = x(x+4)
x(x+4)=4+(x+x+4)
x^2+4x=2x+8
x^2-2x+4x-8=0
x^2+2x-8=0
x^2+4x-2x-8=0
x(x+4)-2(x+4)=0
(x+4)(x-2)=0
x=-4 OR 2
The numbers are (-4,0) , (2,6)
m.ananth@hotmail.ca